Q. An electron revolves with frequency 6.6 * 10 15 r.p.s. around nucleas in circular orbit of radius 0.53A 0 of hydrogen atom, then magnetic field produced at the centre of orbit is- 1. 0.125T 2. 1.25T 3. 12.5T 4. 125T​

Q.  An electron revolves with frequency 6.6 * 1015 r.p.s. around nucleas in circular orbit of radius 0.53A0 of hydrogen atom, then magnetic field produced at the centre of orbit is-                                            1. 0.125T              2.   1.25T               3.  12.5T              4.    125T​  

Grade:12th pass

2 Answers

deepak
7 years ago
see the current due to an electron revolving with the frequency if 6.6x1015 r.p.s is same as that when 6.6x1015 electrons move around the orbit in a second.  
 
hence the total current in the orbit is 6.6 x 1015 x 1.6 x10-19 = 10.56 x 10-4 A
 
now consedering the orbit to be a circular one,
 
\overrightarrow{B}= \nuI / 2R
 
\overrightarrow{B}= 12.5 T
 
hence option 3 is correct
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem below.
 
The total current in the orbit, I = q/T = f x q = 6.6 x 1015 x 1.6 x10-19 = 10.56 x 10-4 A
now consedering the orbit to be a circular one,
Magnetic field at the center of orbit, B = μ0I/2R
Hence, B = 12.5 T
Hence option 3 is correct
 
Hope it helps,
Thanks and regards,
Kushagra